OpenStudy (anonymous):

Solve the equation: -4(x + 10) - 6 = -3(x - 2)

OpenStudy (anonymous):

-4x-40-6=-3x+6 -4x-46=-3x+6 +4x-6 +4x-6 -52=x

OpenStudy (theviper):

\[\Huge{\color{red}{-4x-46=-3x+6}}\]\[\LARGE{\color{green}{\implies-4x+3x=6+46}}\]\[\LARGE{\color{green}{\implies-x=52}}\]\[\LARGE{\color{green}{\implies x=-52}}\]

OpenStudy (anonymous):

@TheViper Is there a reason, why you are rewriting everything that I have already written? If I have the wrong answer, by all means correct me.

OpenStudy (theviper):

@Mrbonez321 I am late in writing the ans as I was doing it in colour that the asker can easily understand:)